I need detail solution:
A grandfather gave a total of 1,000,000 cym (one million Uzbek cym) to his five grandchildren. Starting with the youngest, each got 20,000 cym more than the next younger one. In other words, the youngest got one sum, the next got 20,000 cym more, and so on. How much did the youngest grandchild get?

Answer by rfer(16322)

x+(x+20000)+(x+40000)+(x+60000)+(x+80000)=1000000
5x+200000=1000000
5x=1000000-200000
5x=800000
x=800000/5
x=160,000 cym 1st child